OpenStudy (anonymous):

PLEASE CHEMISTRY please please please please please please help me with these two questions I have tried hundreds of different ways and it always is incorrect :( please help answer!!!!!

OpenStudy (anonymous):

please!

OpenStudy (anonymous):

@thomaster please help me?? my comp is about to die im at the library i have to drive home and get the charger i will be back on though after and it will be within minutes of being due ahhhhhhhhhhhh brb help pleaseeee

thomaster (thomaster):

Sorry i'm not very good with the pressure stuff

OpenStudy (aaronq):

you're basically taking both of the partial pressures and taking both containers as one.

OpenStudy (aaronq):

for the second you need to write an equation and balance it, then using the stoichiometric coefficients, build a relationship in terms of moles. like this general example: aA + bB -> cC + dD lower case = coefficients upper case = species \[\frac{ n _{A} }{ a }=\frac{ n _{B} }{ b }=\frac{ n _{C} }{ c }=\frac{ n _{D} }{ d }\] n means moles, in case you weren't aware.